# Reasoning Efficiently Through Adaptive Chain-of-Thought Compression: A Self-Optimizing Framework

### v1

@@ Abstract.
The introduction of Chain-of-Thought (CoT) reasoning has significantly enhanced the problem-solving capabilities of Large Language Models (LLMs) by prompting them to articulate intermediate reasoning steps rather than producing answers directly. This structured approach has led to significant improvements in accuracy and robustness across benchmarks in arithmetic, logic, and commonsense reasoning. Although effective, it introduces computational overhead. Each reasoning step requires additional token generation, which raises inference latency, memory usage, and key-value cache demands. These issues are particularly critical in domains such as software engineering, where deterministic logic and concise outputs are preferred.
To understand how CoT length impacts both model performance and user experience, we conduct an empirical study on widely used code generation datasets. We find that longer CoT does not always improve performance. Across benchmarks, excessively long reasoning often results in truncation, reduced accuracy, and latency increases of up to five times, with failed outputs consistently longer than successful ones. These findings challenge the assumption that longer reasoning is inherently beneficial and underscore the need for adaptive CoT control. Motivated by these findings, we propose SEER (Self-Enhancing Efficient Reasoning), an adaptive and self-enhancing framework for CoT compression that improves the conciseness of reasoning while preserving output quality, without relying on external compression tools. SEER follows an adaptive self-enhancing paradigm, where CoT responses are refined by combining Best-of-N sampling with task-aware adaptive filtering. The adaptive filter dynamically adjusts thresholds based on the token length of pre-inference model outputs, effectively controlling verbosity while reducing computational costs and preserving accuracy. We evaluate SEER on three software engineering tasks and one mathematical reasoning task. On average, SEER reduces CoT length by 42.1% without compromising performance, and accuracy even improves due to reduced truncation. Moreover, it substantially lowers truncation rates and eliminates most infinite reasoning loops, thereby enhancing both efficiency and robustness. These results highlight SEER as a practical approach to making the use of CoT more efficient, enabling CoT-enhanced LLMs to operate effectively even in resource-constrained settings and on complex tasks.
Chain-of-Thought (CoT), Large Language Models (LLMs), CoT Compression, Software Engineering, Reasoning Efficiency
@@ 1. Introduction
The introduction of Chain-of-Thought (CoT) marks a pivotal advancement in enhancing the reasoning capabilities of Large Language Models (LLMs), particularly for complex, multi-step tasks. By encouraging models to explicitly articulate intermediate reasoning steps, CoT enables a systematic, step-by-step breakdown of problems rather than merely producing an answer in a single attempt (Kojima et al., 2022; Wei et al., 2022). While CoT has demonstrated notable improvements in accuracy and robustness across tasks involving arithmetic, logic, and commonsense reasoning (Nye et al., 2021; Wang et al., 2023; Zhou et al., 2022), its effectiveness becomes even more pronounced in software engineering tasks. Such tasks typically demand precise logical reasoning, structured multi-step problem solving, and transparent decision-making, all of which are well supported by CoT. Recent work shows that CoT can substantially improve performance in multiple software engineering tasks (e.g., code generation, program synthesis, and automated debugging) by improving the interpretability of the model’s reasoning and aligning it with human problem-solving processes (Yao et al., 2023; Li et al., 2025).
However, this benefit comes with substantial costs. As most contemporary LLMs are based on the Transformer architecture (Vaswani et al., 2017), the attention mechanism incurs quadratic computational complexity with respect to context length, while the KV cache grows linearly with context size. Consequently, the verbose decoding of CoT significantly increases inference latency and memory usage, with KV-cache requirements emerging as a key bottleneck (Agrawal et al., 2024; Ho et al., 2024). This inefficiency is particularly critical in software engineering tasks, where timely responses are essential for developer productivity and effective decision-making. Moreover, standard CoT often “overthinks” even simple code generation tasks, generating excessive reasoning tokens despite the existence of straightforward solutions. In more complex scenarios (e.g., repairing software vulnerabilities), this overthinking can impair performance as it increases the inference costs and latencies (Cuadron et al., 2025). Consequently, reducing the length of CoT reasoning without compromising model performance remains a key challenge.
Recent research indicates that compressing the CoT in training data and subsequently fine-tuning the model can effectively manage the length of the generated output while still maintaining accuracy (Kang et al., 2025; Munkhbat et al., 2025a; Xia et al., 2025). Methods such as TokenSkip allow LLMs to eliminate semantically redundant tokens, achieving significant reductions (e.g., from 313 to 181 tokens on the GSM-8K dataset) with minimal performance loss (Xia et al., 2025). Advanced frameworks like Adaptive GoGI-Skip, R1-Compress, and CoLaR have further improved these gains by reducing CoT lengths by 45%–80% while preserving accuracy, as well as delivering 1.6 to 2 times speedups in inference (Zhuang et al., 2025; Wang et al., 2025b; Tan et al., 2025).
However, most of these methods are evaluated on mathematical reasoning datasets and commonsense QA, without assessing their effectiveness on software engineering problems; consequently, they may achieve good performance on mathematical reasoning tasks but exhibit substantially weaker results in software engineering contexts. Moreover, most of these approaches use external compression tools to compress CoT, such as GPT-4 in C3oT (Kang et al., 2025) and LLMLingua in TokenSkip (Xia et al., 2025), which may lead to information loss. Such information loss may lead to “thought leap” and ultimately influence the performance of the model (Xu et al., 2025b). For software engineering tasks, these compression tools may change the logic of the code or even make the code unable to execute.
In this work, we conduct an empirical study to investigate the impact of excessively long CoT outputs on coding tasks using four open-source reasoning models. Our empirical study shows that excessively long CoT reasoning can harm the performance of LLMs. On HumanEval, small LLMs suffer substantial accuracy drops from truncation (e.g., 60.36% to 46.34% on the 1.5B model), while larger models show limited gains since failed outputs are often linked to longer reasoning. In addition, lengthy CoT raises latency by up to five times, underscoring the need to control reasoning length. Motivated by these findings, we propose SEER (Self-Enhancing Efficient Reasoning), an adaptive and self-enhancing framework for CoT compression that improves the conciseness of reasoning while preserving output quality, without relying on external compression tools.
SEER adopts a self-enhancing paradigm, in which the model generates its own CoT-augmented responses and refines them via a two-stage data refinement process: 1) Best-of-N (BoN) sampling, which selects the shortest correct reasoning trace among multiple candidates to ensure both correctness and conciseness, and 2) an adaptive CoT filtering mechanism that dynamically calibrates maximum reasoning length based on the statistical properties of each dataset. This design enables task-specific control over verbosity, mitigating over-compression risks while eliminating pathological repetition patterns. Notably, SEER fine-tuned for one domain (e.g., code generation) also demonstrates strong transferability, preserving concise reasoning in other domains without retraining.
We evaluate SEER on three diverse software engineering tasks and one mathematical reasoning benchmark under an 8K token context constraint. Experimental results show that SEER achieves an average CoT length reduction of 42.1% while preserving or improving accuracy across all tasks. Compared to baselines such as TokenSkip and Naive BoN, our method remains robust in complex software engineering tasks, where other approaches often fail due to instability or logical corruption. Moreover, SEER mitigates infinite reasoning loops, which occur when a model repeatedly generates the same fragment until the context limit is reached, by up to 97.7%, substantially reducing truncation rates. In summary, our main contributions are as follows:
  • We conduct an empirical study on HumanEval to examine how excessively long CoTs affect coding tasks. The results show that overlong reasoning often causes truncation, loops, and latency, ultimately reducing accuracy and stability, motivating the need for effective CoT length control.
@@
  • To address this, we propose SEER, a lightweight CoT compression framework that learns concise reasoning patterns from self-generated outputs, effectively reducing truncation and infinite loops without external modules.
@@
  • Extensive experiments across software engineering and mathematical reasoning tasks demonstrate that SEER achieves an average 42.1% reduction in CoT length while maintaining or even improving pass@1 accuracy.
  • Ablation and cross-domain studies further confirm that adaptive CoT filtering is the key driver of efficiency and accuracy, with BoN sampling providing complementary benefits, and that SEER generalizes well across different task domains.
@@ 2.1. Chain-of-Thought Efficiency
However, the verbosity of CoT introduces significant computational overhead. Each reasoning step requires additional token generation, directly increasing inference latency and resource consumption, particularly for the KV-cache (Agrawal et al., 2024; Ho et al., 2024). This overhead poses a critical bottleneck for real-time applications or systems with constrained computational budgets. Consequently, a key research direction in CoT efficiency is to strike an optimal balance: retaining the reasoning benefits of CoT while minimizing its associated costs. Approaches in this area typically explore methods such as pruning redundant steps, learning to generate more concise reasoning paths (Kang et al., 2025; Xia et al., 2025), or selectively applying CoT based on task complexity (Wang et al., 2025a).
2.2. Overthinking Phenomenon in CoT
@@
Beyond simple inefficiency, the relationship between CoT length and accuracy is not a straightforward “longer is better.” Prior work shows that longer reasoning does not always translate into higher accuracy (Xie et al., 2025). Instead, many studies identify an “overthinking” phenomenon. Accuracy often improves with additional reasoning at the early stages, since the model has more opportunity to explore possible solution paths. However, once the reasoning exceeds an optimal range, performance begins to decline as the model produces redundant or even erroneous content (Chen et al., 2024b; Jin et al., 2024; Cuadron et al., 2025). This decline is a direct result of overthinking, where excessive reasoning introduces noise, deviates from the correct reasoning path, and ultimately undermines accuracy (Chen et al., 2024a). Taken together, these findings highlight that the key challenge is to carefully control CoT length so as to reduce overthinking while preserving accuracy.
2.3. Self-Generated Data for Self-Supervised Fine-Tuning
Recent research has demonstrated that LLMs can bootstrap their own capabilities by generating candidate data, applying automated quality filtering, and then fine-tuning on the filtered subset. STaR illustrates this approach. The model generates CoT rationales and answers, discards incorrect outputs, and fine-tunes iteratively on the correct rationale–answer pairs, reaching reasoning performance comparable to models that are thirty times larger (Zelikman et al., 2022). Additionally, extensions such as Quiet-STaR generalize STaR by generating token-level rationales during continued pretraining, improving zero-shot reasoning on benchmarks like GSM8K and CommonsenseQA without explicit task fine-tuning (Zelikman et al., 2024). These methods collectively form a cohesive pipeline: self-generation, automatic filtering, and self-fine-tuning. This structure allows LLMs to independently improve their reasoning, domain knowledge, and generalization capabilities without the need for human-labeled data.
@@ 3. Empirical Study
Before introducing our proposed framework, we conduct an empirical study to investigate how the length of CoT reasoning influences the performance, efficiency, and stability of LLMs in software engineering scenarios. Our analysis primarily targets coding tasks, a core application area of software engineering where precise and executable reasoning is essential. This study provides a quantitative assessment of the challenges introduced by excessively long CoT outputs, thereby motivating the design of our method.
@@ 3.1.1. Datasets
Our empirical study is conducted on the HumanEval (et al., 2021), which is a widely used benchmark consisting of manually curated Python programming problems. Each problem contains a function signature, a docstring, and a set of test cases for evaluating functional correctness. Additionally, in the final part of our empirical study, we use a more complex dataset, Codeforces (Penedo et al., 2025). Codeforces comprises challenging algorithmic optimization problems collected from one of the most popular websites among competitive programmers.
@@ 3.1.2. LLMs
We employ the DeepSeek-R1-Distill-Qwen (DeepSeek-AI, 2025a) series of models to conduct our empirical study. This choice is motivated by the fact that it is one of the most popular reasoning models in recent times. Moreover, it is open-source, which facilitates deployment. We conduct experiments on four LLMs with sizes of 1.5B, 7B, 14B, and 32B.
@@ 3.1.3. Evaluation Metrics
To assess the performance of CoT, we employ two metrics.

### v2

@@ Abstract.
Chain-of-Thought (CoT) prompting can substantially improve the reasoning ability of large language models (LLMs), but it often comes with high inference cost due to long and poorly controlled reasoning traces. This overhead is particularly problematic in software engineering tasks (e.g., code generation), where both latency and output reliability matter.
To better understand this trade-off, we conduct an empirical study on widely used code generation benchmarks and observe that many modern reasoning models produce excessively verbose CoTs (often thousands of tokens), which frequently leads to truncation and unstable generation. Using a strict $n$-gram repetition detector, we find that the vast majority of truncations are associated with degenerate looping behaviors. Moreover, longer CoT does not necessarily yield better outcomes: failed generations tend to be longer than successful ones, indicating diminishing or even negative returns from overlong reasoning.
Motivated by these findings, we propose SEER (Self-Enhancing Efficient Reasoning), a self-enhancing framework for adaptive CoT compression. SEER improves the conciseness of reasoning while preserving output quality, without relying on external compression tools. SEER refines self-generated CoT data via Best-of-N sampling to suppress looping and redundant traces, then applies a lightweight, data-driven filter to encourage concise yet correct reasoning. It then fine-tunes the model on the filtered data to internalize concise reasoning behaviors. Across three software engineering tasks, SEER reduces CoT length by 41.6% on average while improving task performance, largely by reducing truncation and mitigating reasoning loops.
Chain-of-Thought Compression, Software Engineering, Large Language Models, Reasoning Efficiency
@@ 1. Introduction
The introduction of Chain-of-Thought (CoT) marks a pivotal advancement in enhancing the reasoning capabilities of Large Language Models (LLMs), particularly for complex, multi-step tasks. By encouraging models to explicitly articulate intermediate reasoning steps, CoT enables a step-by-step breakdown of problems rather than producing an answer in a single attempt (Kojima et al., 2022; Wei et al., 2022). Existing studies have shown that CoT achieves notable improvements across tasks, including arithmetic, logic, and commonsense reasoning (Nye et al., 2021; Wang et al., 2023; Zhou et al., 2022). Recently, the utilization of CoT has become more important in software engineering tasks, since they typically require precise logical reasoning, structured multi-step problem solving, and decision-making (Yao et al., 2023; Li et al., 2025).
Figure 1. Examples of Common Chain-of-Thought Issues Across Different Models
Figure 1 presents examples of common CoT issues observed across different open-source reasoning models. Even state-of-the-art reasoning LLMs may generate excessive or poorly controlled CoT, including overthinking, repetitive reasoning, and semantically unproductive deliberation. This issue can arise even in simple tasks with concise solutions. Such verbosity increases inference cost and latency, which is particularly problematic in code‑related scenarios. In software‑engineering contexts where reasoning is part of agent‑based workflows, uncontrolled CoT can increase system overhead and cause unnecessary reasoning steps, as observed in AutoGPT and LangGraph (Significant Gravitas, 2023a; LangChain AI, 2024; Significant Gravitas, 2023b). A common case is that an agent repeatedly re‑examines the same action or planning step without advancing, which wastes tokens and slows execution.
To examine the effect of verbose CoTs in real-world settings, we conduct a systematic empirical analysis of CoT reasoning across a diverse set of open-source models on code generation tasks. We observe that excessively long CoT is common: most reasoning models generate 2K to 4K CoT tokens on average, and truncation can reach up to 17.1% under a 16K token budget. This verbosity is closely associated with reduced efficiency and reliability, as enabling CoT can increase token cost by up to an order of magnitude. Using a strict $n$-gram repetition detector, we find that 236 out of 261 truncations (90.4%) are identified as looping, and manual inspection suggests that this figure is a conservative lower bound. More importantly, our empirical results show that prompt-based conciseness control is limited and highly model-dependent. Overall, these findings highlight the need for effective mechanisms to regulate reasoning length and mitigate looping behaviors, i.e., CoT compression methods.
Prior work on CoT compression can be generally divided into two categories, one of which applies explicit compression operations to CoT, for example, TokenSkip with LLMLingua and C3oT with GPT-4 (Xia et al., 2025; Kang et al., 2025; Zhuang et al., 2025). These methods usually achieve high compression ratios, but they often cause information loss and “thought leap” effects (Xu et al., 2025a). The second line of work relies on repeated sampling or distillation to collect simpler reasoning examples and encourage the model to produce more concise reasoning, such as Naive BoN (Munkhbat et al., 2025; Yu et al., 2024). This form of implicit compression avoids information loss, but typically cannot achieve a high level of compression.
In this paper, we propose SEER (Self-Enhancing Efficient Reasoning), an adaptive framework designed to make CoT reasoning both effective and efficient for software engineering tasks. Rather than relying on external compression tools, SEER internalizes CoT control into the training process by learning concise reasoning patterns directly from self-generated outputs. SEER adopts a self-enhancing paradigm in which the model generates multiple CoT-augmented candidates and refines them through two complementary mechanisms. First, Best-of-N (BoN) sampling explicitly addresses looping behaviors by selecting the shortest correct reasoning trace among multiple candidates, effectively suppressing loops and redundant expansions observed in our empirical study. Second, SEER introduces an adaptive CoT filtering strategy that calibrates maximum reasoning length based on dataset-specific length distributions. This mechanism is grounded in the empirical observation that effective reasoning converges to a narrow length range across models, and prevents both over-compression and excessive verbosity.
We evaluate SEER on three representative software engineering tasks: code generation, defect detection and natural language code search. Across all tasks, SEER reduces CoT length by an average of 41.6% while preserving or even improving pass@1 accuracy. Compared with existing compression baselines, SEER demonstrates superior robustness, achieving the highest pass@1 accuracy while also delivering the most substantial reduction in CoT length on code-related tasks. Importantly, SEER substantially mitigates truncation and infinite reasoning loops. As a result, reasoning loops are mitigated by up to 96.8%, leading to improved inference efficiency and output stability. In summary, our main contributions are as follows:
  • We present a systematic empirical study to analyze the impact of long CoT reasoning in code generation. Our results indicate that excessive CoT is common and truncation is an important factor affecting model performance. Most truncations are associated with reasoning loops, which lower accuracy, raise latency, and make generation unstable.
@@
  • We propose SEER, a lightweight and self-enhancing CoT compression framework that makes CoT reasoning both effective and efficient for software engineering tasks. By learning concise reasoning patterns from self-generated outputs, SEER effectively suppresses reasoning loops and truncation without relying on external compression tools or manual annotations.
@@
  • We evaluate SEER across multiple software engineering tasks. Experimental results show that SEER achieves an average 41.6% reduction in CoT length while preserving or even improving pass@1 accuracy, largely due to improved robustness against truncation and overlong reasoning.
@@ 2.1. The Efficiency of CoT
However, the verbosity of CoT introduces significant computational overhead. Each reasoning step requires additional token generation, directly increasing inference latency and resource consumption, particularly for the KV-cache (Agrawal et al., 2024; Ho et al., 2024). This overhead poses a critical bottleneck for real-time applications or systems with constrained computational budgets. Consequently, a key research direction in CoT efficiency is to strike an optimal balance: retaining the reasoning benefits of CoT while minimizing its associated costs. Approaches in this area typically explore methods such as pruning redundant steps, learning to generate more concise reasoning paths (Kang et al., 2025; Xia et al., 2025), or selectively applying CoT based on task complexity (Wang et al., 2025).
2.2. Existing Issues in CoT
@@
Overthinking Phenomenon in CoT. Beyond simple inefficiency, the relationship between CoT length and accuracy is not a straightforward “longer is better.” Prior work shows that longer reasoning does not always translate into higher accuracy (Xie et al., 2025). Instead, many studies identify an “overthinking” phenomenon. Accuracy often improves with additional reasoning at the early stages, since the model has more opportunity to explore possible solution paths. However, once the reasoning exceeds an optimal range, performance begins to decline as the model produces redundant or even erroneous content (Chen et al., 2024b; Jin et al., 2024; Cuadron et al., 2025). This decline is a direct result of overthinking, where excessive reasoning introduces noise, deviates from the correct reasoning path, and ultimately undermines accuracy (Chen et al., 2024a). Overall, these findings highlight that the key challenge is to carefully control CoT length to reduce overthinking while preserving accuracy.
Looping Problem in CoT. Recent studies have revealed looping problems in CoT reasoning (Pipis et al., 2025), where LLMs repeatedly generate identical or semantically redundant reasoning segments instead of making progress toward a solution. This phenomenon also causes practical issues in real-world applications, as users have reported failures induced by looping behavior on various developer forums and agent frameworks (Significant Gravitas, 2023a; LangChain AI, 2024; Significant Gravitas, 2023b). Looping is particularly prominent under greedy or low-temperature decoding and has been observed across a wide range of modern reasoning models. Prior work suggests that this behavior stems from misestimated action probabilities during training: when advancing to the correct next reasoning step is difficult, models may favor repetitive continuations that are easier to predict, resulting in self-reinforcing loops. Such errors are further amplified by deterministic decoding, allowing repetition to persist across multiple steps. Although increasing sampling randomness can mitigate looping to some extent, it does not address the underlying learning deficiencies, indicating that looping is a fundamental inefficiency of CoT reasoning rather than a purely inference-time artifact.
@@ 3. Empirical Study
We first conduct an empirical study to investigate how the length of CoT reasoning influences the performance, efficiency, and stability of LLMs in software engineering scenarios. We choose a representative problem in software engineering, i.e., code generation tasks, as the subject of our empirical study. This study provides a quantitative assessment of the challenges introduced by excessively long CoT outputs, thereby motivating the design of our method.
@@ 3.1.1. Datasets
Our empirical study is conducted on two widely used benchmarks, i.e., HumanEval (et al., 2021) and MBPP-Sanitized (Austin et al., 2021). They consist of Python programming problems, each accompanied by a function signature, a natural-language docstring describing the task, and a set of unit tests used to evaluate functional correctness..
@@ 3.1.2. LLMs
In our empirical study, we evaluate CoT behavior using a set of open-source reasoning-capable LLMs that are tested under the same experimental settings. The models cover different parameter scales and architectures, including the DeepSeek-R1-Distill-Qwen family (7B, 14B, and 32B), DeepSeek-R1-Distill-Llama-8B, as well as other widely used reasoning models such as QwQ-32B, Qwen3-8B, and gpt-oss-20b.
@@ 3.1.3. Evaluation Metrics
To assess the performance of CoT on software engineering tasks, we employ three metrics.
Pass@1. Pass@1 calculates the proportion of problems for which the model produces a correct solution on the first attempt, directly reflecting its practical effectiveness in solving tasks. It is widely used on code generation tasks as it emphasizes single-attempt correctness under realistic deployment scenarios (et al., 2021).
Average Token Length. We report the average number of tokens generated in model outputs. It reflects the verbosity of the reasoning process and helps assess whether excessively long outputs introduce redundancy or cause truncation. It also facilitates a direct comparison of output lengths among models with and without CoT reasoning.
Truncation Number. We count the number of instances where generation is terminated due to exceeding the maximum token budget. This metric reflects the risk of uncontrolled or redundant reasoning, which may prevent the model from completing valid solutions and thus degrade real-world usability.
3.1.4. Implementation Details
To ensure a fair and reliable empirical evaluation, we describe the implementation details of our experimental setup in this subsection.
Disabling CoT. In our experiments, we evaluate models with explicit CoT disabled. Following the practice adopted by Qwen3-8B (Team, 2025), we suppress intermediate reasoning by appending an empty thinking segment, i.e., “<think> </think>”, to the input prompt. This operation prevents the model from emitting explicit reasoning traces while preserving its ability to produce final answers.
Context Length and Prompt Control. To ensure fair comparison, the maximum context length is fixed to 16k tokens for all models across all experiments. To avoid prompt-induced variation in reasoning length, we exclude any instructions that explicitly encourage or constrain reasoning behavior (e.g., “Please reason step by step”).
For example, for the HumanEval dataset, we use the following instruction:
Please write a python program to address the following QUESTION. Your ANSWER should be in a code block format like this: ```python # Write your code here ```.
Loop Detection. Following Pipis et al. (2025), we identify looping behavior using an $n$-gram repetition criterion. A response is classified as looping if any $n$-gram appears at least $k$ times within the same output. Unless otherwise specified, we set $n=30$ and $k=20$ for all reasoning models, consistent with the original study. This criterion captures severe degenerative repetition patterns commonly observed under greedy or low-temperature decoding and is empirically robust to moderate variations of $n$ and $k$.
3.2. Empirical Findings
How does CoT perform? Table 1 shows the performance of CoT in different models. The results reveal that excessive CoT length is a common phenomenon across most evaluated models. Except for gpt-oss-20b, all models generate long reasoning traces, with average CoT lengths ranging from 2K to 4K tokens. Importantly, non-negligible truncation rates are observed on both HumanEval and MBPP-Sanitized, reaching up to 17.1% in some settings. The results suggest that excessive CoT verbosity is a common issue in modern reasoning models. It appears consistently across diverse architectures and model sizes, frequently exceeding generation budgets.
Table 1. Performance Comparison on HumanEval and MBPP-Sanitized (Max Generated Tokens = 16,384)
Models HumanEval (164 Questions) MBPP-Sanitized (378 Questions)
$pass@1(\%)$ CoT / Total Truncation # $pass@1(\%)$ CoT / Total Truncation #
Qwen3-8B (w/o CoT) 84.76% - / 249 0 (0.0%) 82.01% - / 78 0 (0.0%)
Qwen3-8B 89.02% 3,604 / 3,701 11 (6.7%) 84.13% 3,663 / 3,712 43 (11.4%)
DeepSeek-Qwen-7B (w/o CoT) 65.24% - / 412 0 (0.0%) 50.79% - / 173 0 (0.0%)
DeepSeek-Qwen-7B 78.04% 4,099 / 4,487 28 (17.1%) 73.54% 2,459 / 2,806 33 (8.7%)
DeepSeek-Qwen-14B (w/o CoT) 82.32% - / 456 0 (0.0%) 74.87% - / 424 0 (0.0%)
DeepSeek-Qwen-14B 87.20% 3,282 / 3,669 18 (11.0%) 86.51% 2,078 / 2,453 17 (4.5%)
DeepSeek-Qwen-32B (w/o CoT) 87.20% - / 183 0 (0.0%) 88.62% - / 199 0 (0.0%)
DeepSeek-Qwen-32B 88.41% 3,136 / 3,539 17 (10.4%) 88.10% 2,333 / 2,686 28 (7.4%)
DeepSeek-Llama-8B 76.22% 3,640 / 4,021 19 (11.6%) 63.23% 2,568 / 2,910 35 (9.3%)
QwQ-32B 93.90% 2,363 / 2,458 2 (1.2%) 93.65% 1,929 / 1,990 6 (1.6%)
gpt-oss-20b 92.68% 481 / 790 1 (0.6%) 94.44% 470 / 693 3 (0.8%)
    DeepSeek-Llama-8B, QwQ-32B and gpt-oss-20b do not support disabling CoT.
To evaluate the impact of CoT, we compare the results with and without CoT on models that support disabling CoT reasoning. Enabling CoT improves pass@1 accuracy by an average of 7.4% across the evaluated models, indicating that CoT effectively enhances reasoning performance. Specifically, the performance gains are substantially larger for smaller models (e.g., +12.8% on HumanEval for DeepSeek-Qwen-7B), indicating that CoT plays a more critical role when model capacity is limited.
However, these accuracy improvements come at a substantial efficiency cost. Enabling CoT increases the token cost by up to an order of magnitude in some cases (e.g., from approximately 400 tokens to over 4K tokens for DeepSeek-Qwen-7B), leading to significantly higher inference latency. Moreover, the resulting long CoT frequently triggers truncation, which partially offsets the accuracy benefits of CoT. In several settings, truncated CoT leads to degraded performance compared to non-CoT generations, particularly when the final answer is cut off. For example, on MBPP-Sanitized, DeepSeek-Qwen-32B shows a small decrease in pass@1 (from 88.62% to 88.10%), alongside a truncation rate of 7.4%.
Overall, these results highlight a fundamental trade-off in the use of CoT. While CoT improves reasoning accuracy, its effectiveness is constrained by excessive verbosity, increased latency, and truncation-induced failures. This observation motivates a closer look at why CoT becomes excessively long and how much of the generated reasoning is truly needed for solving the problem, which we explore in the next sections.
    Finding 1:  CoT improves accuracy but induces excessive verbosity and truncation. Across various models, enabling CoT improves pass@1 accuracy by an average of 7.4%. However, it can increase token cost about 10 times and frequently cause truncation, with a maximum rate of 17.1%.
How Common is Looping among Truncations? To quantify how often truncation is associated with looping behavior, we analyze all truncated outputs using the loop detection criterion described in Section 3.1.4. Across all 261 truncation cases observed across different models and datasets, 236 cases (90.4%) are identified as looping by the $n$-gram repetition criterion.
Since our loop detector is intentionally strict (n=30n{=}30, k=20k{=}20), borderline cases with weaker or shorter repetition may be missed. We manually examine the remaining truncation cases not flagged by the detector and find that some of them still exhibit loop-like repetition patterns, although they do not satisfy the strict $n$-gram threshold. Therefore, the 91.8% should be viewed as a conservative lower bound, and the true fraction of truncations affected by looping is likely higher.
    Finding 2:  Looping is highly common in truncated outputs. Using a strict $n$-gram repetition detector, 91.8% of truncations are identified as looping, and manual inspection suggests this is a conservative estimate.
How Much Reasoning is Effective? To quantitatively assess the proportion of useful reasoning within CoT, we conduct a LLM-based CoT compression. Following prior work that employs strong LLMs as judges or analyzers to evaluate model-generated text (Zheng et al., 2023), we use GPT-4o (OpenAI, 2024) to extract the effective reasoning steps in CoT outputs. Specifically, this process removes the surface-level verbosity while preserving the original logic. Essential reasoning steps and planning instructions are retained. The prompt used for this section is provided in the replication package. To reduce randomness in LLM-based compression, we perform the compression three times for each CoT and report the averaged results.
After compression, we measure 1) the token length of the compressed CoT and 2) the effective reasoning ratio, defined as the fraction of tokens retained after compression relative to the original CoT length. The results are reported in Table 2. Truncated instances are omitted because they have not finished the reasoning process.
Table 2. Effective Reasoning Length and Retention Ratio Estimated via GPT-4o on HumanEval and MBPP
Models HumanEval (164 Questions) MBPP-Sanitized (378 Questions)
Avg. CoT (Before) Avg. CoT (After) Retained (%) Avg. CoT (Before) Avg. CoT (After) Retained (%)
DeepSeek-Llama-8B 3,594 212 5.91 2,538 164 6.50
DeepSeek-Qwen-7B 3,677 210 5.71 2,334 177 7.62
DeepSeek-Qwen-14B 3,076 211 6.88 1,974 181 9.17
DeepSeek-Qwen-32B 2,869 212 7.42 2,218 179 8.07
QwQ-32B 2,311 216 9.36 1,866 187 10.05
Qwen3-8B 3,456 228 6.62 3,527 155 4.41
gpt-oss-20b 481 149 31.11 470 128 27.27
Table 2 shows significant redundancy in generated CoTs. For most models (except gpt-oss-20b), the compressed lengths consistently fall between 150 and 230 tokens across benchmarks. Consequently, the effective reasoning ratio falls below 10%, suggesting that most generated tokens are unnecessary for the core reasoning. Additionally, gpt-oss-20b shows more efficient reasoning, but approximately 70% of its reasoning is still ineffective.
A closer comparison between Qwen3-8B and QwQ-32B on MBPP-Sanitized further illustrates this phenomenon. Although Qwen3-8B generates nearly twice as many CoT tokens as QwQ-32B before compression, its effective reasoning ratio after compression is even lower. This indicates that increased CoT verbosity does not correspond to more effective reasoning and may instead amplify redundant or low-value content. Overall, these results indicate that longer CoT is not necessary, and motivate the need for CoT compression that focuses on preserving essential reasoning steps.
    Finding 3:  Only a small fraction of CoT contributes to essential reasoning. Compression-based analysis with GPT-4o shows that, despite large variation in raw CoT length, less than 10% of CoT tokens are retained across most models. The results indicate substantial redundancy in generated reasoning.
Is Longer Reasoning Always Better? To evaluate whether longer CoT reasoning leads to better outcomes, we generate multiple responses for the same problem and analyze the CoT lengths of passed and failed solutions. We focus on a relatively difficult problem, where the quality of CoT is more likely to influence the final result. Since HumanEval does not provide an official difficulty classification, we select the problem with the longest prompt in the HumanEval dataset, i.e., HumanEval/129. We choose the DeepSeek-Qwen-Series Model for our evaluation, focusing on how the length of CoT impacts its performance across various model sizes. The prompt follows the implementation details of this section (Section 3.1.4), without any explicit instructions to control the length of the CoT. For each of the models, we generate 10,000 responses and examine the distribution of CoT lengths, as shown in Figure 2, with the maximum token length set to 16,384.
While prior work (Jin et al., 2024) suggests that increasing the length of CoT reasoning improves accuracy, our findings reveal a different pattern across all three models. Failed outputs are frequently associated with longer CoT lengths, while successful outputs are more concentrated around shorter lengths. For instance, in the 7B model, the median CoT length of failed generations (9,489 tokens) surpasses that of successful generations (8,296 tokens), with a similar pattern observed in both 14B and 32B models. A closer examination of failure cases shows that overly long reasoning often introduces redundancy, propagates errors, and may lead to truncation, which ultimately degrades performance. These results highlight that effective reasoning does not simply require “more thinking”, but rather appropriate reasoning lengths that balance sufficiency and conciseness.
    Finding 4:  Longer CoT does not guarantee higher accuracy and is often associated with failures. On HumanEval/129, failed cases consistently show longer CoT than successful ones across all models (e.g., 9,489 vs. 8,296 tokens in 7B), indicating that excessive reasoning is correlated with failures.
Can Prompt Design Effectively Reduce CoT Length? Prior work has shown that the verbosity of CoT reasoning is sensitive to instruction-level guidance (Wei et al., 2022; Wang et al., 2023; Agrawal et al., 2024). Therefore, to verify whether the prompt design can effectively shorten the length of the CoT, we include a lightweight prompting baseline to encourage concise reasoning. Specifically, we augment the original task instruction with a general conciseness guideline, asking the model to provide its reasoning as concisely as possible while avoiding redundant or unnecessary steps. Concretely, we add the following system prompt:
Please provide your reasoning as concisely as possible, avoiding redundant or unnecessary steps.
Table 3. Effect of Conciseness-Guided Prompting on CoT Length and Model Accuracy
Models HumanEval (164 Questions) MBPP-Sanitized (378 Questions)